Circular No. 7883

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 H2 R. Huber, Lincoln Laboratory, Massachusetts Institute of Technology,

reports the discovery by LINEAR of a comet with a diffuse coma that is elongated in p.a. ∼ 250^{o}. Following posting on the NEO Confirmation Page, R. Apitzsch (Wildberg, Germany; 0.24-m reflector) reported the object to be cometary, and L. Šarounová (Ondřejov; 0.65-m reflector) found a coma of diameter > 1′, m_{2} = 15.5, and a faint fan tail apparently toward the southwest in a crowded star field. CCD astrometry:

2002 UT α_{2000} δ_{2000} m_{1} Observer
Apr. 22.41264 19^{h}29^{m}47.^{s}00 +28^{o}11′00′′.0 15.9 LINEAR
22.44106 19 29 44.50 +28 13 13.7 15.9  _{”}
22.98194 19 28 58.85 +28 55 27.6 16.2 Apitzsch
23.04289 19 28 53.70 +29 00 15.3 15.5  _{”}
23.04964 19 28 53.06 +29 00 45.8  Šarounová
23.05593 19 28 52.50 +29 01 15.6 13.5  _{”}

COMET C/2002 C1 (IKEYA-ZHANG) H. E. Matthews, National Research Council of Canada and Joint As-

tronomy Centre (JAC), Hilo; T. B. Lowe, JAC; and D. Jewitt, Institute for Astronomy, University of Hawaii, report sub-millimeter continuum observations of C/2002 C1 made at the James Clerk Maxwell Telescope (+ SCUBA bolometer array receiver) on Mauna Kea. Images of a 2′ field were made simultaneously at 850 and 450 μm on Mar. 30.85 UT. Preliminary peak intensities are 0.4 and 2.0 Jy/beam respectively, within the beamwidths of 13′′.5 and 8′′.5 (to half-power). The equivalent blackbody cross-section at 850 μm is ≈ 3200 km^{2}, ∼ 8 times that measured for 1P/Halley but 10 times less than that of C/1995 O1 when at its brightest. The 850-μm image is significantly extended compared with that of Uranus and CRL 2688, proving that most of the emission is from large particles in an extended coma surrounding the nucleus. Matthews et al. estimate that the mass loss in these large particles is on the order of 4 × 10^{8} g/s.

Naked-eye total-magnitude estimates: Mar. 29.80 UT, 3.3 (M. Meyer, Ilhorn, Germany); Apr. 6.84, 3.6 (N. Biver, Domart, France; 5^{o} tail); 11.38, 3.4 (P. Creed, Cambridge, OH; 8^{o} tail in p.a. 340^{o}); 14.12, 4.3 (J. Serant, Chevillon, France); 18.6, 3.9 (S. J. O’Meara, Volcano, HI; 6^{o} tail in p.a. 300^{o}); 20.15, 4.2 (J. J. Gonzalez, Asturias, Spain); 22.02, 3.9 (K. Hornoch, Lelekovice, Czech Republic).

2002 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green